OpenStudy (anonymous):

1722+253+278+447=? please help me out guys

OpenStudy (lxelle):

1722+253 = 1975 1975+ 278 = 2253 2253+ 447 =?

OpenStudy (anonymous):

2700

OpenStudy (rational):

1722+253+278+447 (1722+278) + (253+447) (1700+200+100) + (200+400+100) (2000) + (700) (2700)

OpenStudy (beccaboo333):

I can't tell if you really needed help or are just that bored and posted a random question.

OpenStudy (anonymous):

http://lmgtfy.com/?q=1722%2B253%2B278%2B447%3D

OpenStudy (rational):

you need to group them in a way that allows u to calculate it mentally without using pen/paper/google/calculator/others

OpenStudy (lxelle):

Yes you are right.

OpenStudy (anonymous):

thanks guys a lot

OpenStudy (anonymous):

how it takes u to do such operation ? @rational

OpenStudy (anonymous):

mentally :O

OpenStudy (rational):

the problem is designed to be worked mentally

OpenStudy (anonymous):

ik >.< im just asking cuz im sort of slow

OpenStudy (rational):

if I ask you to calculate : 99 + 101 would you use calculator ?

OpenStudy (rational):

how about 111 + 89 ?

OpenStudy (anonymous):

sure i would not :P but see ur very smart in noting this >.< i only note it when u posted ( just asking ) i dont have that talent to see it quick